# Prediction intervals for overdispersed binomial endpoints and their application to historical control data

### v2

@@ 3.2 Bootstrap calibrated prediction intervals
A similar algorithm was used by Menssen et al. 2024 in order to calibrate prediction intervals for overdispersed count data. Both algorithms are modified versions of the bootstrap calibration procedure of Menssen and Schaarschmdit 2022. The only difference to Menssen and Schaarschmdit 2022 is, that both interval limits are calibrated individually, but the search for $q^{calib}_{l}$ and $q^{calib}_{u}$ depends on the same bisection procedure using a tolerance $t=0.001$.
@@
$\Big{[}l=n^{*}\hat{\pi}-q^{calib}_{l}\sqrt{\frac{\hat{\phi}n^{*2}\hat{\pi}(1-\hat{\pi})}{\sum_{h}n_{h}}+\hat{\phi}n^{*}\hat{\pi}(1-\hat{\pi})}$
$\quad\quad\quad u=n^{*}\hat{\pi}+q^{calib}_{u}\sqrt{\frac{\hat{\phi}n^{*2}\hat{\pi}(1-\hat{\pi})}{\sum_{h}n_{h}}+\hat{\phi}n^{*}\hat{\pi}(1-\hat{\pi})}\Big{]}.$
@@
$\Big{[}l=n^{*}\hat{\pi}-q^{calib}_{l}\sqrt{n^{*}\hat{\pi}(1-\hat{\pi})[1+(n^{*}-1)\hat{\rho}]+[\frac{n^{*2}\hat{\pi}(1-\hat{\pi})}{\sum_{h}n_{h}}+\frac{\sum_{h}n_{h}-1}{\sum_{h}n_{h}}n^{*2}\hat{\pi}(1-\hat{\pi})\hat{\rho}]}$
$\quad\quad\quad u=n^{*}\hat{\pi}+q^{calib}_{u}\sqrt{n^{*}\hat{\pi}(1-\hat{\pi})[1+(n^{*}-1)\hat{\rho}]+[\frac{n^{*2}\hat{\pi}(1-\hat{\pi})}{\sum_{h}n_{h}}+\frac{\sum_{h}n_{h}-1}{\sum_{h}n_{h}}n^{*2}\hat{\pi}(1-\hat{\pi})\hat{\rho}]}\Big{]}.$
@@ 3.3 Bayesian modeling
As above, a situation with historical experiments $h=1,...,H$ is considered, but individual observations $i=1,...,I$ per experiment are possible. Let $n_{hi}$ be the total numbers of experimental units and $Y_{hi}\in\{0,1,...,n_{hi}\}$ the numbers of successes among them. As before, the data are modeled on experiment level by aggregation:
$n_{h}=\sum_{i=1}^{I}n_{hi}$
$Y_{h}=\sum_{i=1}^{I}Y_{hi}$
Furthermore, for Bayesian modeling the total number of experimental units is assumed to be constant over all clusters (historical and current control groups):
$n_{hi}=:n^{\prime}\quad\Rightarrow\quad n_{h}=n^{*}=In=:n\quad\forall\ h.$
@@ 3.3.1 Hierarchical modeling
If we assume the binomial model
$Y_{hi}\sim\text{Bin}(n_{hi},\pi_{h})=\text{Bin}(n^{\prime},\pi_{h})$
for individual values, it follows
$Y_{h}\sim\text{Bin}(n,\pi_{h})$
per historical control group.
This is the same situation as above, and again a Beta-binomial model is applied to represent the between-experiment-overdispersion. In a Bayesian interpretation, the experiments’ parameters $\pi_{h}$ are assumed to follow a Beta distribution, as a second level in a hierarchy (Howley and Gibbert 2003).
A Beta prior in mean-precision-parametrization (Beta proportion distribution) is applied (Stan development Team 2024a):
$\pi_{h}\sim\text{Be}(\mu,\kappa)\quad\forall\;h$
@@
A posterior predictive distribution of a future observation $y^{*}$ is obtained by random sampling per MCMC iteration $c=1,...,C$: For every MCMC sample $(\hat{\mu}_{c},\hat{\kappa}_{c})$ of $\mu$ and $\kappa$, a prediction
$\hat{\pi}^{*}_{c}\sim\text{Be}(\hat{\mu}_{c},\hat{\kappa}_{c})$
@@
$y^{*}_{c}\sim\text{Bin}(n,\hat{\pi}^{*}_{c}).$
A pointwise prediction interval for one future observation $y^{*}$ is then obtained using empirical quantiles of $\left\{y^{*}_{1},...,y^{*}_{C}\right\}$ such that
$[l,u]=[q_{\alpha/2}(\left\{y^{*}_{1},...,y^{*}_{C}\right\}),q_{1-\alpha/2}(\left\{y^{*}_{1},...,y^{*}_{C}\right\})]$
@@ 3.3.2 Bayesian generalized linear mixed model
A concept closely related to the above Beta-binomial model is a generalized linear mixed model (GLMM), where Bernoulli distributed random variables $Z_{h1},...,Z_{hn}$ with $\sum_{k=1}^{n}Z_{hk}=Y_{h}$ are considered (Fong et al. 2010). The success proportion $\pi_{h}=E(Z_{hk})$ is linked to a linear predictor
@@
which is comprised of a fixed general intercept $\nu$ and random intercepts $\beta_{h}$ per experiment:
@@
The posterior distributions of $\nu\in(-\infty,+\infty)$ and $\sigma>0$ are estimated by MCMC sampling using non-informative vague priors. Realizations of the future random variable $Y^{*}=\sum_{k=1}^{n}Z^{*}_{k}$ are derived similarly to above: For every MCMC sample $c$, a prediction
$\hat{\beta}^{*}_{c}\sim N(0,\hat{\sigma}_{c})$
@@
$\hat{\eta}^{*}_{c}=\hat{\nu}_{c}+\hat{\beta}^{*}_{c}$
and the future experiment’s success proportion
$\hat{\pi}^{*}_{c}=\frac{\exp(\hat{\eta}^{*}_{c})}{\exp(\hat{\eta}^{*}_{c})+1}$
are derived, and in turn
$y^{*}_{c}\sim\text{Bin}(n,\hat{\pi}^{*}_{c})$
is drawn from the latter. Similarly as above, a pointwise two-sided prediction interval for one future observation $y^{*}$ is given by
$[l,u]=[q_{\alpha/2}(\left\{y^{*}_{1},...,y^{*}_{C}\right\}),q_{1-\alpha/2}(\left\{y^{*}_{1},...,y^{*}_{C}\right\})].$
@@ 3.4 Computational details and estimation
The proposed frequentist prediction intervals are implemented in the R package predint (Menssen 2023). The uncalibrated prediction intervals (eq. 10 and 11) are provided via qb_pi() and bb_pi(). The bootstrap calibrated prediction intervals (eq. 3.2 and 3.2) are implemented via the functions quasi_bin_pi() and beta_bin_pi().
The estimates $\hat{\pi}$ and $\hat{\phi}$ used for the quasi-binomial prediction interval were estimated based on a generalized linear model (stats::glm()). The parameter estimates for the beta-binomial distribution $\hat{\pi}$ and $\hat{\rho}$ were estimated following the approach of Lui et al. 2000. Both estimates that define overdispersion were restricted to avoid underdispersion. The quasi-binomial interval was computed with $\hat{\phi}\geq 1.001$. The estimated intra-class correlation used for the beta-binomial interval was restricted to $\hat{\rho}\geq 0.00001$.
The bisection procedure used for bootstrap calibration is implemented in predint and provided via the function bisection() This function takes three different lists as input, that contain the bootstrapped expected observations $\hat{y}^{*}_{b}$, the bootstrapped standard errors $\widehat{se}(\hat{y}^{*}-Y^{*})_{b}=\sqrt{\widehat{var}({n^{*}\hat{\pi})_{b}+\widehat{var}(Y^{*})_{b}}}$ and the bootstrapped further observations $y^{*}_{b}$ as input and returns values for $q_{l}$ and $q_{u}$. Hence, this function enables the calculation of bootstrap calibrated prediction intervals in a general way, such that any Wald-type prediction interval can be calibrated, as long as a parametric model can be fit to the data from which parametric bootstrap samples can be sampled.
The Bayesian hierarchical beta-binomial model was applied using the software stan via the R package rstan (Stan Development Team 2024). A weakly informative prior $\kappa\sim\text{Ga}(2,\>5\cdot 10^{-5})$ was used, giving the domain of the precision hyperparameter $\kappa>0$. The Bayesian GLMM was applied using the R function stan_glmer from the package rstanarm (Goodrich et al. 2024). Functions for the computation of both Bayesian prediction intervals are provided via GitHub https://github.com/MaxMenssen/pi_overdisp_binomial/tree/main.
@@ 4.1 HCD from the micro-nucleus test
The occurrence of a micronucleus is an adverse event during cell division, where a small fragment of chromosomes remains outside the new nuclei. Hence, an increase of cells with a micronucleus can be considered as a proxy for the genotoxic properties of a chemical compound (Fenech 2000).
In the in-vitro micro-nucleus test (MNT), cell suspensions are cultivated on a well plate of which some are treated with the test compounds (treatment groups), some are untreated (negative control) and some are treated with a compound with known genotoxic potency to evaluate the test assay’s functionality (positive control). At the end of the study, it is examined how many cells (out of a total number of cells) have developed at least one micronucleus.
The setting of the following simulations are inspired by a real life HC data base comprised of negative and positive control groups from about 50 experiments, divided in three parts of about 15 to 20 experiments according to different compounds used for the positive controls. Every control group consists of six wells and $3000$ cells were screened each, such that the aggregated number of cells with at least one micronucleus out of a total of $18000$ cells was considered.
This HC data base hints at an average proportion of cells with at least one micronucleus of about $\hat{\pi}\approx 0.01$ for negative and about $\hat{\pi}\approx 0.15$ to $0.2$ for positive controls. The overdispersion was estimated to range between $\hat{\phi}\approx 10$ and $50$ for negative controls and tend to reach several hundreds for positive controls.
@@ 4.2 HCD from long-term carcinogenicity studies
Historical control data on different endpoints obtained in long-term carcinogenicity studies (LTC) are provided via the Historical Control Database of the National Toxicology Program (NTP 2024). HCD about the number of B6C3F1 mice per control group that had a hemangioma, the number of mice that showed at least one malignant tumor as well as the number of mice that passed away before the end of the study was analyzed by Menssen and Schaarschmidt 2019 with regard to the binomial proportion $\hat{\pi}$ and the amount of overdispersion $\hat{\phi}$.
For rare events (hemangioma) the reported proportion was relatively low (close to zero) and overdispersion was estimated to be absent in most of the cases. The proportion of animals with malignant tumors ranged between 0.2 and 0.8, whereas the estimated dispersion parameters raised up to roughly four, meaning that the particular data set is four times as variable than possible under simple binomial distribution. The reported mortality rates behaved in a similar fashion.
@@ 5 Simulation study
In order to asses the coverage probabilities of the different methods for the calculation of HCL reviewed above, Monte-Carlo simulations were run. In all simulations, the nominal coverage probability was set to $0.951-\alpha=$ 0.95. The parameter combinations used for the simulation were inspired by the real life data shown above. For all simulations the cluster size was fixed ($n_{h}=n^{*}\quad\forall\quad h=1,\ldots,H$) such that the variance-mean relationship of the beta-binomial distribution and the quasi-binomial assumption is not in contradiction to each other.
@@ 5.1 Coverage probabilities for the MNT-inspired simulation setting
In order to assess the coverage probabilities in a setting that mimics the properties of the real life HCD from the micro-nucleus test, the coverage probabilities for the following 72 combinations of $H={5,10,20,100}$ (the number of historical control groups), $\pi={0.001,0.01,0.1}$ (the binomial proportion) and $\phi={1.001,3,5,10,50,500}$ (the amount of overdispersion) was assessed for each of the six methods (three heuristics and three prediction intervals). Since cluster size was relatively huge ($n_{h}=n^{*}=18000$) most parameter combinations result in a relatively symmetrical underlying distribution (except for $\phi=500$, see supplementary fig. 1).
It is obvious, that the historical range (fig. 1A) covers the future observation $y^{*}_{s}$ in all the cases, if the amount of historical control groups is high enough (say above 20). Therefore, the historical range should not be applied for the computation of HCL.
If overdispersion is practically absent ($\phi=1.001$), the np-chart (fig. 1B) yields coverage probabilities relatively close to the nominal level. With increasing overdispersion (and hence, increasing right-skeweness), the coverage probabilities of the np-chart decrease below 0.2. This demonstrates, that the np-chart should not be applied to toxicological (overdispersed) binomial data.
@@
At first glance, it seems that the HCL computed by the mean $\pm$ 2 SD (fig 1C) yield coverage probabilities close to the nominal level of 0.95, if the amount of historical control groups is high enough (at least 20). But, with a decreasing binomial proportion $\pi$ and/or an increasing amount of overdispersion ($\phi$) the right-skeweness of the data increases. Hence, in many cases the lower border almost always covers the future observation (the triangles approach 1) whereas the coverage probability of the upper border (crosses) does not approach the desired 0.975. This behavior demonstrates that HCL computed based on the mean $\pm$ 2 SD do not ensure for equal tail probabilities, or in other words: They do not properly cover the central 95 % of the underlying distribution, if this distribution is skewed.
Figure 1: Simulated coverage probabilities of heuristical methods (MNT-setting) A: Historical range, B: np-chart, C: Mean $\pm$ 2 SD, Black horizontal line: Nominal coverage probability $\psi^{cp}=0.95$, Grey horizontal line: Nominal coverage probability for the lower and the upper limit, if equal tail probabilities are achieved $\psi^{l}=\psi^{u}=0.975$. Note the different scales of the y-axis of the sub-graphics (0.6-1 for A and C, 0-1 for B).
The coverage probabilities of the three different prediction intervals are depicted in Figure 2. For most of the settings ($\pi\geq 0.01$ and $\phi\in[3,50]$), the beta-binomial and the quasi-binomial prediction intervals (Fig. 2 A and B) behave in a comparable manner and approach the nominal coverage probability of 0.95, even if only five historical control groups are available. In this settings, the Bayesian prediction interval (Fig. 2C) approaches the nominal coverage probability only for cases where the number of historical control groups is high ($H>20$), but remains liberal for a lower number of historical control groups.
@@
In the (practical) absence of overdispersion ($\phi=1.01$), the beta-binomial prediction interval controls the alpha error best, whereas the quasi-binomial prediction interval tends to behave conservatively. But, with an increasing amount of historical control groups, also this interval approaches the nominal coverage probability. Contrary to this, the prediction interval obtained from Bayesian hierarchical modeling remains conservative in this setting, even for 100 historical studies.
@@
For moderate proportions $\pi\in[0.001,01]$ and moderate to high overdispersion $\phi\in[3,50]$ (moderate in the context of $n=18000$), the quasi- and beta-binomial prediction intervals yield coverage probabilities close to the nominal 0.95, whereas the interval obtained from Bayesian hierarchical modeling approaches the nominal level only for 100 historical studies and remains liberal for a lower amount of historical information.
@@
The right-skewness of the underlying distribution increases with a decreasing binomial proportion and an increasing amount of overdispersion (panels in the lower left of the three sub-graphics). In this case, the sampled data contains many zeros and hence, the computed lower borders always cover the future observation (the triangles approach 1). Hence, the prediction intervals practically aim to yield 97.5 % upper prediction borders. This effect is clearly visible for all three intervals and reflects a core feature of the data rather than being based on model misspecification.
@@
Figure 2: Simulated coverage probabilities of prediction intervals (MNT-setting) A: Calibrated beta-binomial prediction interval, B: Calibrated quasi-binomial prediction interval, C: Prediction interval obtained from a Bayesian hierarchical model, Black horizontal line: Nominal coverage probability $\psi^{cp}=0.95$, Grey horizontal line: Nominal coverage probability for the lower and the upper limit, if equal tail probabilities are achieved $\psi^{l}=\psi^{u}=0.975$

### v3

@@ 3.2 Bootstrap calibrated prediction intervals
[…]
@@
[…]
The applied bootstrap calibration is a modified version of the algorithm proposed by Menssen and Schaarschmdit 2022 and has also been successfully applied to overdispersed count data (Menssen et al. 2024). The only difference to Menssen and Schaarschmdit 2022 is, that both interval limits are calibrated individually, but the search for $q^{calib}_{l}$ and $q^{calib}_{u}$ depends on the same bisection procedure using a tolerance $t=0.001$.
Furthermore, the algorithm presented above differs from the calibration procedure used by Menssen and Schaarschmidt 2019 in two important aspects: First, contrary to Menssen and Schaarschmidt 2019 it calibrates both prediction borders separately and hence accounts for equal tail probabilities (see fig. 4 of the supplementary materials). Second, it aims to find substitutes for the complete standard normal quantile used for interval calculation, which in principal also enables simultaneous prediction of the outcome of more than one control group. This is possible if $y^{*}_{b}$ is substituted by a bootstrap sample that mimics the outcome of several control (or treatment) groups.
@@ 3.3.1 Hierarchical modeling
Bayesian hierarchical modeling depends on the beta-binomial model that is given in equations 1 and 2. In a Bayesian interpretation, the experiments’ parameters $\pi_{h}$ are assumed to follow a Beta distribution, as a second level in a hierarchy (Howley and Gibbert 2003).
A Beta prior in mean-precision-parametrization (Beta proportion distribution, Stan development Team 2024a) is applied:
$\pi_{h}\sim\text{Beta}(\mu,\kappa)\quad\forall\;h$
@@
A posterior predictive distribution of a future observation $y^{*}$ is obtained by random sampling per MCMC iteration $c=1,...,C$: For every MCMC sample $(\tilde{\mu}_{c},\tilde{\kappa}_{c})$ of $\mu$ and $\kappa$, a prediction
$\tilde{\pi}^{*}_{c}\sim\text{Beta}(\tilde{\mu}_{c},\tilde{\kappa}_{c})$
@@
$\tilde{y}^{*}_{c}\sim\text{Bin}(n^{*},\tilde{\pi}^{*}_{c}).$
A pointwise prediction interval for one future observation $y^{*}$ is then obtained using empirical quantiles of $\left\{\tilde{y}^{*}_{1},...,\tilde{y}^{*}_{C}\right\}$ such that
$[l,u]=[q_{\alpha/2}(\left\{\tilde{y}^{*}_{1},...,\tilde{y}^{*}_{C}\right\}),q_{1-\alpha/2}(\left\{\tilde{y}^{*}_{1},...,\tilde{y}^{*}_{C}\right\})]$
@@ 3.3.2 Bayesian generalized linear mixed model
A concept closely related to the Beta-binomial model shown above is a generalized linear mixed model (GLMM), where Bernoulli distributed random variables $Z_{h1},...,Z_{hn_{h}}$ with $\sum_{k=1}^{n_{h}}Z_{hk}=Y_{h}$ are considered (Fong et al. 2010). The success proportion $\pi_{h}=E(Z_{hk})$ is linked to a linear predictor
@@
which is comprised of a fixed general intercept $\nu$ and random effects $\beta_{h}$ per experiment:
@@
The posterior distributions of $\nu\in(-\infty,+\infty)$ and $\sigma>0$ are estimated by MCMC sampling using non-informative vague priors. Realizations of the future random variable $Y^{*}=\sum_{k=1}^{n^{*}}Z^{*}_{k}$ are derived: For every MCMC sample $c$, a prediction
$\tilde{\beta}^{*}_{c}\sim N(0,\tilde{\sigma}_{c})$
@@
$\tilde{\eta}^{*}_{c}=\tilde{\nu}_{c}+\tilde{\beta}^{*}_{c}.$
Then, the future experiment’s success proportion
$\tilde{\pi}^{*}_{c}=\frac{\exp(\tilde{\eta}^{*}_{c})}{\exp(\tilde{\eta}^{*}_{c})+1}$
is derived, and in turn
$\tilde{y}^{*}_{c}\sim\text{Bin}(n^{*},\tilde{\pi}^{*}_{c})$
is drawn from the latter. Similarly as above, a pointwise two-sided prediction interval for one future observation $\tilde{y}^{*}$ is given by
$[l,u]=[q_{\alpha/2}(\left\{\tilde{y}^{*}_{1},...,\tilde{y}^{*}_{C}\right\}),q_{1-\alpha/2}(\left\{\tilde{y}^{*}_{1},...,\tilde{y}^{*}_{C}\right\})].$
@@ 3.4 Computational details and estimation
The proposed frequentist prediction intervals are implemented in the R package predint (Menssen 2023). The uncalibrated prediction intervals (eq. 11 and 12) are provided via qb_pi() and bb_pi(). The bootstrap calibrated prediction intervals (eq. 13 and 14 ) are implemented via the functions quasi_bin_pi() and beta_bin_pi().
The estimates $\hat{\pi}$ and $\hat{\phi}$ used for the quasi-binomial prediction interval were estimated based on a generalized linear model (stats::glm()). The estimates $\hat{\pi}$ and $\hat{\rho}$ for the beta-binomial distribution were obtained following Lui et al. 2000. In order to avoid undersipersion, the quasi-binomial interval was computed with $\hat{\phi}\geq 1.001$. The estimated intra-class correlation used for the beta-binomial interval was restricted to $\hat{\rho}\geq 0.00001$.
The bisection procedure used for bootstrap calibration is implemented in predint and provided via the function bisection(). This function takes three different lists as input, that contain the bootstrapped expected observations $\hat{y}^{*}_{b}$, the bootstrapped standard errors $\sqrt{\widehat{var}({n^{*}\hat{\pi})_{b}+\widehat{var}(Y^{*})_{b}}}$ and the bootstrapped further observations $y^{*}_{b}$ and returns values for $q_{l}$ and $q_{u}$. Hence, this function enables the calculation of bootstrap calibrated prediction intervals in a general way, such that any Wald-type prediction interval can be calibrated, as long as a parametric model can be fit to the data and a sampling function is available based on which parametric bootstrap samples can be generated.
The Bayesian hierarchical beta-binomial model was applied using the software stan via the R package rstan (Stan Development Team 2024). In the simulation, a weakly informative prior $\kappa\sim\text{Ga}(2,\>5\cdot 10^{-5})$ was used, giving the domain of the precision hyperparameter $\kappa>0$. The Bayesian GLMM was applied using the R function stan_glmer from the package rstanarm (Goodrich et al. 2024). Functions for the computation of both Bayesian prediction intervals are provided via GitHub https://github.com/MaxMenssen/pi_overdisp_binomial/tree/main.
@@ 4 Properties of real life HCD
This section reviews the statistical properties of two real life HC data bases: One about long-term carcinogenicity studies and one about the micronucleus test. The properties of the two data bases were taken into account tor the setup of the Monte-Carlo simulations regarding the coverage probability of the different HCL. This two study types were chosen in order to reflect the different properties of overdispersed binomial data at very different locations in the possible parameter space in order to enhance the simulations generalizability.
@@ 4.1 HCD from the micro-nucleus test
A micronucleus is a small fragment of chromosomes remaining outside the new nuclei after cell division. In the in-vitro micro-nucleus test (MNT), cell cultures on a well plate are treated with the compound of interest and an increased number of cells with a micronucleus is used as a proxy for genotoxicity (Fenech 2000). Negative and positive control wells are observed to measure the spontaneous micronucleus proportion without treatment and to evaluate the test assay’s functionality by comparison to HCD.
The analyzed real life HCD includes negative and positive control groups from about $H\approx 50$ experiments. Due to the high extend of standardization, the aggregated number of scored cells per control group is always $n_{h}=n^{*}=18000$.
The HC data base hints at an average proportion of cells with at least one micronucleus of about $\hat{\pi}\approx 0.01$ for negative control groups and about $\hat{\pi}\approx 0.15$ to $0.2$ for positive controls. Overdispersion was estimated to range between $\hat{\phi}\approx 10$ and $50$ for negative controls and tends to reach several hundreds for positive controls.
@@ 4.2 HCD from long-term carcinogenicity studies
Historical control data on different endpoints obtained in long-term carcinogenicity studies (LTC) are publicly available via the Historical Control Database of the U.S. National Toxicology Program (NTP 2024). From this resource Menssen and Schaarschmidt 2019 extracted HCD about the numbers of B6C3F1 mice per control group that had a hemangioma, the numbers of mice that showed at least one malignant tumor as well as the numbers of mice that passed away before the end of the study. This data was analyzed with regard to the binomial proportion $\hat{\pi}$ and the amount of overdispersion $\hat{\phi}$.
For rare events (hemangioma) the reported proportion was relatively low (close to zero) and overdispersion was estimated to be absent in most of the cases. The proportion of animals with malignant tumors ranged between 0.2 and 0.8, whereas the estimated dispersion parameters raised up to roughly four, meaning that the particular data set is four times as variable than possible under simple binomial distribution. The reported mortality rates behaved in a similar fashion. All reported control groups were comprised of 50 animals.
@@ 5 Simulation study
In order to asses the coverage probabilities of the different methods for the calculation of HCL reviewed above, Monte-Carlo simulations were run. In all simulations, the nominal coverage probability was set to $0.951-\alpha=$ 0.95. The parameter combinations used for the simulation were inspired by the real life data shown above. For all simulations the cluster size was fixed ($n_{h}=n^{*}\quad\forall\quad h=1,\ldots,H$) such that the variance-mean relationships of the beta-binomial distribution and the quasi-binomial assumption are not in contradiction to each other.
@@
Average interval borders were calculated as the mean of the simulated interval borders
$\bar{l}=\frac{\sum_{s=1}^{S}l_{s}}{S}$
$\bar{u}=\frac{\sum_{s=1}^{S}u_{s}}{S}.$
@@ 5.1 Coverage probabilities for the MNT-inspired simulation setting
This part of the simulation was inspired by the properties of the real live HCD for the micro-nucleus test reviewed above. 72 combinations of values for the number of historical studies $H$, the binomial proportion $\pi$ and the amount of overdispersion $\phi$ were run, of which 18 combinations directly reflect the properties of the real life HCD (combinations of the bold faced parameter values in tab. 1), whereas the remaining 54 combinations of parameter values were run in order to enhance the generalisability of the simulation.
Note that, because the cluster size was relatively huge ($n_{h}=n^{*}=18000$) most parameter combinations result in a relatively symmetrical underlying distribution (except for small proportions $\pi=0.001$ combined with high overdispersion $\phi=500$, see supplementary fig. 1).
Table 1: Parameter values used for the MNT-inspired simulation
Parameter Values1
No. of historical studies ($H$) 5, 10, 20, 100
Binomial proportion ($\pi$) 0.001, 0.01, 0.1
Overdispersion ($\phi$)2 1.001, 3, 5, 10, 50, 500
Historical cluster size ($n_{h}$) 18000
Future cluster size ($n^{*}$) 18000
Bold numbers: Parameter values that reflect estimates obtained from real life HCD. 1: Simulations were run for all 72 combinations of the given parameter values. 2: Since the cluster sizes $n_{h}$ and $n\textdegree*$ are constant $\phi\widehat{=}1+(n_{h}-1)\rho\widehat{=}1+(n^{*}-1)\rho$. The given values of $\phi$ correspond to intra-class correlation coefficients $\rho$ of 5.5e-08, 0.00011, 0.00022, 0.00050, 0.00272, 0.02772.
The simulated coverage probabilities of the heuristical methods are depicted in fig. 2, whereas the coverage probabilities of the proposed prediction intervals are presented in fig. 3. Furthermore, simulated average interval borders and the corresponding true quantiles are given in fig. 4 for three simulation settings, that reflect different levels of possible skewness of the underlying distribution. In these simulation settings H was set to 100 historical control groups. Thus, the historical information is high enough, that obvious flaws of the methods should be clearly visible and not be caused by too low historical sample size. None of the heuristics control the central 95 % of the underlying distribution, whereas all prediction intervals properly converge to the true quantiles, regardless if the underlying distribution is skewed or not.
As expected, the historical range (fig. 2A) covers the simulated future observations $y^{*}_{s}$ in all the cases, if the amount of historical control groups is high enough and lacks a proper control of the statistical error (here 5 %).
If overdispersion is practically absent ($\phi=1.001$), the control limits of the np-chart (fig. 2B) yield coverage probabilities relatively close to the nominal level.But, with increasing overdispersion, their coverage probabilities decrease below 0.2, meaning that, by far, less than the desired central 95 % of the underlying distribution are covered. This is also demonstrated in fig. 4: The average limits of the np-chart are covered by the true quantiles. Consequently, the np-chart can not be recommended for the application to toxicological (overdispersed) HCD.
@@
At first glance, it seems that the HCL computed by the mean $\pm$ 2 SD (fig. 2C) yield coverage probabilities close to the nominal level of 0.95, if the amount of historical control groups is high enough (at least 20). But, with a decreasing binomial proportion $\pi$ and/or an increasing amount of overdispersion $\phi$ the right-skewness of the data increases. Hence, in many cases the lower border almost always covers the future observation (the triangles in fig. 2C approach 1) whereas the coverage probability of the upper border (crosses) does not approach the desired 0.975. This behavior demonstrates that HCL computed based on the mean $\pm$ 2 SD do not ensure for equal tail probabilities, or in other words: They do not control the central 95 % of the underlying distribution, if this distribution is skewed. This behavior is also demonstrated in fig. 4: The simulated average limits are systematically below the corresponding true quantiles, if the underlying distribution is right-skewed.
The coverage probabilities of the prediction intervals proposed in sections 3.2 and 3.3 are depicted in Figure 3. In general, both frequentist prediction intervals and the interval based on a Bayesian GLMM yield coverage probabilities relatively close to the nominal level for most settings. All three intervals behave in a comparable fashion, except for settings in which the number of historical control groups is low and / or the amount of overdispersion is high ($H<10$, $\phi=500$). Contrary to this, the prediction interval drawn from a Bayesian hierarchical model, fails to control the statistical error in many of the cases. Furthermore, Bayesian hierarchical models faced serious convergence issues, if the amount of historical studies and the amount of overdispersion was high ($H=100$, $\phi=500$), meaning that the desired interval could not be derived for many of the simulated data sets. Consequently, the coverage probability could not be computed in this three cases (which explains the missing dots for $H=100$ and $\phi=500$ in 3D).
@@
In the (practical) absence of overdispersion ($\phi=1.01$), the beta-binomial prediction interval controls the alpha error best, whereas the prediction intervals based on the quasi-binomial assumption or a Bayesian GLMM tend to behave conservatively. But, in the absence of overdispersion, both intervals approach the nominal coverage probability with an increase of historical control groups. Contrary to this, the prediction interval obtained from Bayesian hierarchical modeling remains conservative in this setting, even for 100 historical studies.
@@
For low to moderate proportions $\pi\in[0.001,0.1]$ and low to moderate overdispersion $\phi\in[3,50]$ (moderate in the context of $n_{h}=n^{*}=18000$), all intervals, except the one based on Bayesian hierarchical modeling, yield coverage probabilities close to the nominal 0.95. The interval obtained from Bayesian hierarchical modeling approaches the nominal level only for 100 historical studies and remains liberal for a lower amount of historical information.
@@
The right-skewness of the underlying distribution increases with a decreasing binomial proportion and an increasing amount of overdispersion (panels in the lower left of each sub-graphic). In this case, the sampled data contains many zeros and hence, the computed lower borders tend to always cover the future observation (the triangles approach 1). Hence, the prediction intervals practically aim to yield 97.5 % upper prediction borders. This effect is also visible in fig. 4 and reflects a core feature of the underlying distribution, rather than being based on model misspecification.
@@
Figure 2: Simulated coverage probabilities of heuristical methods (MNT-setting). A: Historical range, B: np-chart, C: Mean $\pm$ 2 SD, Black horizontal line: Nominal coverage probability $\psi^{cp}=0.95$, Grey horizontal line: Nominal coverage probability for the lower and the upper limit, if equal tail probabilities are achieved $\psi^{l}=\psi^{u}=0.975$. Coverage probabilities below 0.6 are excluded from the graphic.
Figure 3: Simulated coverage probabilities of prediction intervals (MNT-setting). A: Calibrated beta-binomial prediction interval, B: Calibrated quasi-binomial prediction interval, C: Prediction interval obtained from a Bayesian generalized linear mixed model, D: Prediction interval obtained from a Bayesian hierarchical model, Black horizontal line: Nominal coverage probability $\psi^{cp}=0.95$, Grey horizontal line: Nominal coverage probability for the lower and the upper limit, if equal tail probabilities are achieved $\psi^{l}=\psi^{u}=0.975$. Coverage probabilities below 0.9 are not shown. The Bayesian hierarchical model faced serious convergence problems for $\phi=500$ and $H=100$. Hence, coverage probabilities could not be computed in this setting.
Figure 4: Average limits vs. true quantiles of the underlying distribution (MNT-setting). Grey area: Probability mass function of the underlying beta binomial distribution. Black lines: True underlying 2.5 % and 97.5% quantiles. Dashed red lines: Average limits for $H=100$ and $n_{h}=18000$ obtained from the simulation.